# HA-VLN: A Benchmark for Human-Aware Navigation in Discrete–Continuous Environments with Dynamic Multi-Human Interactions, Real-World Validation, and an Open Leaderboard

### v1

@@ Abstract
Vision-and-Language Navigation (VLN) systems often focus on either discrete (panoramic) or continuous (free-motion) paradigms alone, overlooking the complexities of human-populated, dynamic environments. We introduce a unified Human-Aware VLN (HA-VLN) benchmark that merges these paradigms under explicit social-awareness constraints. Our contributions include: (1) a standardized task definition that balances discrete–continuous navigation with personal-space requirements; (2) an enhanced human motion dataset (HAPS 2.0) and upgraded simulators capturing realistic multi-human interactions, outdoor contexts, and refined motion–language alignment; (3) extensive benchmarking on 16,844 human-centric instructions, revealing how multi-human dynamics and partial observability pose substantial challenges for leading VLN agents; (4) real-world robot tests validating sim-to-real transfer in crowded indoor spaces; and (5) a public leaderboard supporting transparent comparisons across discrete and continuous tasks. Empirical results show improved navigation success and fewer collisions when social context is integrated, underscoring the need for human-centric design. By releasing all datasets, simulators, agent code, and evaluation tools, we aim to advance safer, more capable, and socially responsible VLN research.***Project: https://ha-vln-project.vercel.app/
Figure 1: HA-VLN Navigation Scenario. An agent navigates environments with dynamic human activities, adjusting its path based on natural language instructions and real-time sensor observations to avoid collisions. Key positions (e.g., , ) align with instructional cues referring to specific human behaviors. When the agent encounters a bystander on the phone (, Decision A), it intelligently turns right to avert a potential collision. On the right, RGB and Depth observations illustrate the agent’s panoramic view preceding decisions A, B, and C, capturing its dynamic responses to nearby humans. [Best viewed in color and high resolution.]
11footnotetext: Equal contribution. Work done during internship at UW.22footnotetext: Corresponding author.
@@ 1 Introduction
Vision-and-Language Navigation (VLN) [3, 56] enables robots to understand multimodal instructions and navigate real or simulated spaces [15, 49]. While recent VLN systems have shown encouraging results in constrained domains, most focus exclusively on either discrete (panoramic) or continuous (free-motion) navigation, omitting the rich human dynamics and partial observability of realistic environments [4, 23, 54]. These omissions compromise real-world applicability, as embodied agents must handle unpredictable human motion, avoid collisions, and process instructions referencing socially salient cues.
@@
Motivation & Gaps. Despite notable progress, three major challenges persist in today’s VLN research. First, social awareness remains underexplored: human participants in the scene are commonly overlooked or reduced to inert obstacles, preventing the agent from respecting personal space or reacting to bystanders’ activities (see Figure 1). Second, instruction complexity is not well captured in existing corpora. Commands such as “Go around the chair” rarely reflect real-world contexts like “Go around the person talking on the phone” [39]. Third, static-environment assumptions dominate, neglecting real-time re-planning when people traverse corridors or gather spontaneously. In practice, social navigation demands partial observability and dynamic route adjustment.
Human-Aware VLN. To tackle these limitations, Human-Aware VLN (HA-VLN) [31] explicitly integrates dynamic human interactions, driving agents to navigate among people engaged in various everyday behaviors. Although HA-VLN moves closer to realistic settings, it remains constrained by four main issues (Appendix Table A1): (1) Limited Human Motion Data, as initial datasets (e.g., HAPS 1.0 [31]) sometimes misalign textual labels and 3D motion sequences; (2) Discrete Navigation Bias, since many simulators restrict navigation to panoramic hops [28]; (3) Underdeveloped Multi-Human Interactions, typically modeling only a single person in indoor scenes; and (4) Instruction Generation Deficits, where automated text expansions omit time-varying human behaviors, limiting agents’ adaptability in dynamic, crowded environments.
Our Approach & Contributions. In this paper, we substantially extend HA-VLN by unifying discrete and continuous navigation, coupling high-fidelity human motion data, multi-agent interactions, and real-time rendering under a single umbrella. Concretely, we make four primary contributions, each elaborated in its respective section:
@@
  • Standardized Task & Metrics [Sec. 2]. We unify discrete and continuous navigation under social-awareness constraints, ensuring consistent goals and evaluations.
@@
  • Upgraded Data & HA-VLN Simulator [Sec. 3]. Building on HAPS 1.0 [31], we present HAPS 2.0 (486 SMPL sequences) and two advanced simulators (HA-VLN-DE, HA-VLN-CE) that incorporate multi-view human annotation, dual-thread rendering, and rigorous collision checks for up to 910 active individuals (Figure 2).
@@
  • Comprehensive Benchmarking [Sec. 4]. We enrich R2R-CE with 16,844 human-centric instructions (Table A3) and benchmark multiple agents under unified metrics, unveiling challenges arising from multi-human dynamics and partial observability.
@@
  • Real-World Validation & Leaderboard [Sec. 5]. We robustly demonstrate sim-to-real transfer using a physical robot successfully navigating crowded indoor areas, and provide a public leaderboard for comprehensive discrete–continuous evaluations in multi-human scenarios.
As visualized in Figures 1 and 2, introducing realistic human motion significantly amplifies task complexity, leading to pronounced performance drops in baseline VLN models. By bridging discrete and continuous navigation, enriching human-motion fidelity, and validating results both in simulation and on physical robots, our approach narrows the gap between conventional VLN setups and real-world, socially aware navigation systems. We release all related datasets, simulators, agent code, and evaluation tools to further fuel progress in this emerging research area.
Figure 2: HA-VLN-CE Simulator. We extend photorealistic Habitat environments by integrating diverse and dynamic human activities using the HAPS 2.0 dataset (172 activities, 486 motion models, 58,320 frames). The annotation pipeline proceeds in two stages: (i) a coarse-to-fine optimization approach with PSO and synchronized multi-view cameras, and (ii) a human-in-the-loop iterative refinement process to enrich multi-human interactions and movements. Real-time rendering uses a signaling mechanism to update human motion dynamically, enabling continuous collision detection and agent-environment interplay. [Best viewed in color and high resolution.]
2 Human-Aware VLN Task
Motivation & Overview. We present Human-Aware Vision-and-Language Navigation (HA-VLN), a next-generation VLN framework that specifically requires agents to navigate among dynamically moving people while following natural-language instructions. Unlike standard VLN—where humans are either absent or treated as static obstacles—HA-VLN compels agents to anticipate human motion, respect personal space, and adapt paths to avoid collisions. As shown in Figure 1, the agent must interpret instructions referencing ongoing activities (e.g., “Go upstairs where someone is pacing on the phone…”) and then maneuver accordingly in a photorealistic 3D space.
State & Action Spaces. At each time step $t$, the agent’s state is
$s_{t}\;=\;\langle\mathbf{p}_{t},\,o_{t},\,\Theta_{t}^{\text{FOV}}\rangle,$
where $\mathbf{p}_{t}$ is its 3D location, $o_{t}$ its orientation, and $\Theta_{t}^{\text{FOV}}$ its egocentric view. In discrete environments (DE), agents hop among pre-defined viewpoints, each providing an RGB observation. In continuous environments (CE), agents see an RGB+D feed with a $90^{\circ}$ field of view and can move in small increments (e.g., $0.25\,\mathrm{m}$ forward, $15^{\circ}$ turn). Across both settings, the action space is
$\mathcal{A}\;=\;\{\,a_{\text{forward}},\,a_{\text{left}},\,a_{\text{right}},\,a_{\text{up}},\,a_{\text{down}},\,a_{\text{stop}}\}.$
Figure 2 provides examples of how these actions evolve within complex, human-occupied scenes.
Human-Aware Constraints. Beyond reaching a goal, agents must obey strict social protocols:
  • Dynamic Human Models. People move naturally according to 3D motion trajectories from HAPS 2.0, which are continuously updated in real time.
  • Personal Space. Agents must avoid getting too close (under 3 m in DE or within the sum of radii in CE).
  • Human-Focused Instructions. Language often describes people and their activities (e.g., “Go around the person talking on the phone”), necessitating alignment between textual prompts and live visual inputs.
These detailed human placements and movements are annotated via a multi-stage pipeline (Sec. 3), entailing extensive labeling and verification to ensure realism.
Dynamic & Partial Observations. Because people can move unpredictably, each time step is a partially observable Markov decision process (POMDP). The new state $s_{t+1}$ depends on both the agent’s action and concurrent human movements (e.g., someone stepping aside or emerging from a corridor). Agents must infer unobserved factors—such as whether a person will yield space—and balance exploration (seeking alternative routes) and exploitation (continuing on a known path) to reach the target efficiently.
Challenges & DE–CE Synergies. Building socially compliant robots in crowded settings poses three key Sim2Real challenges: (i) Social Navigation, i.e., collision-free motion that respects personal space; (ii) Human-Aligned Instructions, where language may refer to fleeting activities or interactions; and (iii) Adaptive Re-Planning, as humans can block or unblock hallways unpredictably. While DE allows rapid prototyping via discrete viewpoint jumps, CE approximates real-world dynamics and motion fidelity. By integrating both approaches, HA-VLN spans the spectrum from large-scale simulation to practical on-robot deployment, significantly broadening research in socially aware, human-centric navigation.
@@ 3 HA-VLN Simulator
Motivation & Overview. The HA-VLN Simulator addresses long-standing challenges in socially-aware navigation by placing multiple, dynamically moving humans into both discrete and continuous 3D environments. While many existing simulators either overlook human behaviors or model them as static obstacles, our simulator features high-fidelity motion, multi-person interactions, and real-world complexities such as group gatherings, spontaneous movements, and personal-space constraints. Table A2 situates our approach relative to Habitat 3.0 [42] and HabiCrowd [48], highlighting broader scenario coverage, more natural motion types, and explicit instruction alignment. Meanwhile, Table A3 contrasts HAPS 2.0 with its predecessor, HAPS 1.0 [31], underscoring notable advancements in motion diversity, alignment accuracy, and overall realism. By building on HAPS 2.0, we leverage 486 motion sequences that span indoor and outdoor activities. We provide two complementary modules: HA-VLN-CE for continuous navigation and HA-VLN-DE for discrete navigation. Both modules share a unified API (Sec. 3), offering consistent access to human states, dynamic scene updates, and collision checks. Figure 2 presents an overview of how these elements converge within an agent’s action-observation loop.
HAPS 2.0 Dataset. Evolving from HAPS 1.0 [31], HAPS 2.0 encompasses 486 3D human-activity descriptions covering 26 region types (e.g., living rooms, gyms, sidewalks). Each description (e.g., “workout gym exercise: An individual running on a treadmill”) is validated through human surveys and ChatGPT-4 [6] for plausibility. We employ the Motion Diffusion Model (MDM) [17] to transform these text descriptions into 120-frame SMPL mesh sequences, $\mathcal{H}=\langle h_{1},\ldots,h_{120}\rangle$, capturing detailed 3D pose and shape information. Figure A2 illustrates various indoor and outdoor contexts, while Figure A3 depicts representative motions (e.g., climbing stairs, running).
Annotation Pipeline: Coarse-to-Fine. To position these motions in large-scale 3D scans, we employ a two-stage annotation procedure. Stage 1 (Coarse) uses Particle Swarm Optimization (PSO) [25] (Algorithm A1) to allocate motions within each region $\mathbf{R}$, bounded by $\mathbf{B}_{\text{lo}},\mathbf{B}_{\text{hi}}$, and link them to suitable objects $\mathbf{O}$. A subset $\mathbf{H}^{\prime}\subseteq\mathbf{H}$ is filtered by region labels, then paired with objects via semantic cues (e.g., “sitting on a chair” aligns with a chair object). PSO penalizes boundary violations and personal-space intrusions (1 m), producing highly plausible initial placements. Stage 2 (Fine) refines these placements with a precise nine-camera approach [22, 40] (Figure A1). Annotators iteratively adjust rotations and positions to eliminate collisions or interpenetrations (e.g., limbs intersecting walls), requiring over 430 hours of human labeling to achieve 529 precisely aligned models across 90 scans. To further enhance diversity, we integrate a human-in-the-loop process [11] for generating multi-person scenarios: large language models propose contextual interactions (Listings LABEL:list:dataset_prompt and LABEL:list:few-shot-example), which annotators then refine. Figure A3 illustrates enhanced interactions (e.g., group discussions). After both stages, we assemble 910 human models distributed over 428 regions in 90 scans, including 268 complex motions (e.g., running, stair climbing). Figure A5 and Appendix B.8 detail their distributions, revealing extensive coverage of diverse motion types and scene layouts.
@@
Figure 3: Motion Analysis. (a) Top three motions from Stage 1 (without enrichment) and Stage 2 (with enrichment). (b) Overall statistics of human activities, comparing interaction types, movement distances, and the number of models used. Enrichment extends both the variety and dynamic aspects of human activities.
Real-Time Rendering & Agent Interaction. Our simulator updates human motions in real time rather than relying on static captures. A dual-thread producer-consumer architecture (Algorithm A2) orchestrates frame updates, with Thread 1 enqueuing “refresh” signals and Thread 2 processing them in sync with the agent’s action cycle. Each motion spans up to 120 frames; once Thread 2 detects a signal, it discards old meshes and loads new ones, maintaining retrieval delays under 50 ms. Figure A2 exemplifies how multiple humans can simultaneously populate a shared environment. Agents perceive these moving entities through a navigation mesh (navmesh) [45]. Collisions are flagged if the distance between bounding volumes is less than their combined radii, leading to an automatic revert. This mechanism ensures agents learn to respect personal space and navigate effectively around dynamic crowds.
@@
Discrete vs. Continuous Settings. HA-VLN-CE (Continuous) allows agents to move in real-valued increments (e.g., $0.25\,\mathrm{m}$ forward, $15^{\circ}$ turns), supporting fine-grained collision avoidance and social reactivity. As Figure A5 shows, each scene can contain up to 10 humans, with simulation speeds reaching 30–60 FPS on standard GPUs. HA-VLN-DE (Discrete) extends HA3D [31] to incorporate HAPS 2.0 data, encompassing both indoor and outdoor settings. Here, agents hop among panoramic viewpoints while humans continue to move, preserving core social-navigation challenges. We map continuous positions to discrete viewpoints [31], add small offsets to refine alignment, and integrate 627 annotated humans across 90 buildings.
@@
Unified API.  We provide a unified API for both modes with three core functionalities: (1) Human State Queries retrieve bounding volumes, motion frames, and semantic annotations of nearby humans; (2) Dynamic Scene Updates notify agents of newly moved humans or environmental changes; and (3) Collision Checks detect whether a proposed move (e.g., forward step or viewpoint hop) would result in intersection with a human. By unifying HAPS 2.0 with coarse-to-fine annotation, real-time multi-human rendering, and a single API across discrete and continuous settings, our HA-VLN Simulator significantly broadens the scope of socially-aware navigation research. Figures A2, A3, and A5 depict the simulator’s capacity to capture rich human behaviors, while Tables A1 and A3 underscore our advantages over earlier simulators, including Habitat 3.0 and HabiCrowd, as well as the improvements of HAPS 2.0 relative to HAPS 1.0. Appendix B.7 outlines environment sizes, latency considerations, and usage examples.
Figure 4: HA-R2R Dataset Analysis. (a) Instruction-length distribution by human group size (no humans to more than three). (b) Comparison of instruction lengths between HA-R2R and R2R-CE.

### v3

# HA-VLN 2.0: An Open Benchmark and Leaderboard for Human-Aware Navigation in Discrete and Continuous Environments with Dynamic Multi-Human Interactions

@@ Abstract
Vision-and-Language Navigation (VLN) has been studied mainly in either discrete or continuous settings, with little attention to dynamic, crowded environments. We present HA-VLN 2.0, a unified benchmark introducing explicit social-awareness constraints. Our contributions are: (i) a standardized task and metrics capturing both goal accuracy and personal-space adherence; (ii) HAPS 2.0 dataset and simulators modeling multi-human interactions, outdoor contexts, and finer language–motion alignment; (iii) benchmarks on 16,844 socially grounded instructions, revealing sharp performance drops of leading agents under human dynamics and partial observability; and (iv) real-world robot experiments validating sim-to-real transfer, with an open leaderboard enabling transparent comparison. Results show that explicit social modeling improves navigation robustness and reduces collisions, underscoring the necessity of human-centric approaches. By releasing datasets, simulators, baselines, and protocols, HA-VLN 2.0 provides a strong foundation for safe, socially responsible navigation research.
11footnotetext: Equal contribution.22footnotetext: Corresponding author (zhiqics@uw.edu).
@@ 1 Introduction
Vision-and-Language Navigation (VLN) Anderson et al. (2018); Zhang et al. (2024b) challenges embodied agents to interpret natural-language instructions and reach specified goals in photorealistic simulators or real-world environments Gu et al. (2022); Wang et al. (2022). Although recent advances have delivered strong performance in controlled benchmarks, existing methods are typically confined to either discrete (DE) or continuous (CE) settings, neglecting the complexities of crowded, human-populated spaces, where agents must contend with unpredictable human behaviors, reason under partial observability, and ensure socially compliant navigation Anderson et al. (2021); Kadian et al. (2020); Yu et al. (2024). Bridging these gaps is essential for moving VLN from simulation prototypes toward robust real-world deployment Wu et al. (2024); Gao et al. (2024).
@@
Motivation and Open Challenges. Despite recent progress, VLN research still faces three fundamental limitations that restrict its real-world applicability. First, social awareness remains underexplored: human participants in the scene are commonly overlooked or reduced to inert obstacles, preventing the agent from respecting personal space or reacting to bystanders’ activities (see Figure 1). Second, finer-grained instructions are not well captured in existing corpora Paduraru et al. (2021); Kong et al. (2024). Commands such as “Turn to your left, and go past the chair” rarely reflect real-world contexts like “Turn to your left, where you will see someone taking a brief pause … on the chair” in Figure 1. Third, static-environment assumptions neglects real-time re-planning when people traverse corridors or gather spontaneously. In practice, social navigation demands partial observability and dynamic route adjustment. Addressing these issues requires a benchmark that unifies DE and CE with explicit regime disclosure, supports socially grounded finer-grained instructions, and incorporates human-centric metrics for navigation in dynamic multi-human environments.
Figure 1: HA-VLN 2.0 Navigation Scenario. HA-VLN 2.0 adds four key challenges: (i) unified discrete/continuous navigation with denser crowds, richer activities, and mixed indoor–outdoor scenes; (ii) stricter social-distance and collision constraints under partial observability; (iii) instructions explicitly grounded in human activities and spatial cues, improving language–vision alignment; and (iv) robust real-time planning amid occlusion and multi-human dynamics. Example: key positions (e.g., , ) align with instructional cues referring to specific human behaviors. When the agent encounters a bystander on the phone (, Decision A), it intelligently turns right to avert a potential collision. On the right, RGB and Depth observations illustrate the agent’s panoramic view preceding decisions A, B, and C, capturing its dynamic responses to nearby humans.
Toward Human-Aware VLN. Early progress, notably HA-VLN 1.0 framework Li et al. (2024) introduced dynamic humans into VLN, yet several shortcomings limited its realism and reproducibility (Appendix Table A1). Motion data in HAPS 1.0 Li et al. (2024) suffered from alignment errors and limited diversity, restricting coverage of everyday activities. The benchmark also exhibited a discrete-navigation bias, with its simulator largely confined to viewpoint hops Krantz et al. (2020) rather than physics-consistent low-level control Krantz et al. (2021). Multi-human interactions were underdeveloped, typically modeling only a single individual in simplified scenarios. Finally, instruction generation remained coarse and object-centric, omitting temporally varying activities and offering little control over granularity. These limitations call for a benchmark that standardizes regime disclosure, expands motion fidelity and diversity, incorporates multi-human interactions, and supports finer-grained socially grounded instructions across both discrete and continuous settings.
Our Contributions. In response, we introduce HA-VLN 2.0, a unified benchmark coupling discrete (DE) and continuous (CE) navigation paradigms with explicit social-awareness constraints. It comprises the HAPS 2.0 dataset, featuring 486 SMPL-based motion sequences across 26 region types, rigorously annotated via multi-view verification (around 430 annotation hours). HA-VLN 2.0 includes established simulators (HA-VLN-DE, HA-VLN-CE) with multi-human interactions, outdoor environments, real-time rendering, and precise collision management for up to 910 human models across 428 regions in 90 scans. A unified API enables seamless comparisons across modes (Fig. 2; Sec. 3). Additionally, we expand R2R-CE Krantz et al. (2020) with 16,844 socially grounded instructions and introduce two robust baseline agents, HA-VLN-VL with Transformer-based grounding and HA-VLN-CMA with cross-modal attention for replanning, both validated under comprehensive human-centric metrics (Sec. 4). Finally, we demonstrate successful sim-to-real transfer in real-world robot validation and provide a public evaluation leaderboard (Sec. 5.2).
Specifically, HA-VLN 2.0 offers four key advancements:
  1. 1.
@@
    Cross-paradigm task standardization & Metrics. We unify DE and CE navigation under social-awareness constraints, ensuring consistent goals and human-centric evaluations (Sec. 2).
  2. 2.
@@
    HAPS 2.0 & Dual simulators (large-scale build). We release HAPS 2.0 (486 SMPL sequences) and two established simulators (HA-VLN-DE, HA-VLN-CE) that incorporate multi-view human annotation ($\sim$ 430 human-hours), outdoor scenes, dual-thread rendering, and rigorous collision checks for up to 910 active individuals with interactions (Fig. 2; Sec. 3).
  3. 3.
@@
    Comprehensive benchmarking with finer-grained instructions. We augment R2R-CE with 16,844 socially-grounded instructions and benchmark multiple agents under unified metrics, unveiling challenges arising from multi-human dynamics and partial observability. (Sec. 4).
  4. 4.
@@
    Real-robot validation and public leaderboard. We robustly demonstrate sim-to-real transfer using a physical robot successfully navigating crowded indoor areas, and provide a public leaderboard for comprehensive discrete–continuous evaluations in multi-human scenarios (Sec. 5.2).
2 The Unified Human-Aware VLN Task
Motivation and Overview. HA-VLN 1.0 Li et al. (2024) introduced dynamic humans into VLN, but its discrete-environment (DE) focus limited ecological validity and hindered systematic study of continuous control and realistic multi-human interactions. To address this, we formalize HA-VLN 2.0, a unified benchmark that integrates DE and CE under explicit human-centric constraints. Under this setting, agents must parse instructions that reference ongoing human activities (e.g., “Go upstairs where someone is pacing on the phone”), anticipate plausible human trajectories, maintain socially compliant distances, and adapt plans online in densely populated, photorealistic 3D scenes (Fig. 1). We next make this specification precise by unifying state and action across regimes.
Unified State and Action Spaces. HA-VLN 2.0 defines a shared state and action interface bridging DE and CE. At each timestep $t$, the agent state is
$s_{t}=\langle\mathbf{p}_{t},\,o_{t},\,\Theta_{t}^{\text{FOV}}\rangle,$ (1)
where $\mathbf{p}_{t}$ is the agent’s 3D position, $o_{t}$ its orientation, and $\Theta_{t}^{\text{FOV}}$ its egocentric visual observation. In DE, agents hop among predefined viewpoints with RGB observations; in CE, they perceive RGB+D within a $90^{\circ}$ field of view and execute fine-grained increments (e.g., $0.25\,\mathrm{m}$ forward, $15^{\circ}$ rotation). Crucially, DE and CE now share a unified action space,
$\mathcal{A}=\{a_{\text{forward}},\,a_{\text{left}},\,a_{\text{right}},\,a_{\text{up}},\,a_{\text{down}},\,a_{\text{stop}}\},$ (2)
enabling direct and fair comparison across paradigms (Fig. 2).
Human-Aware Enhanced Constraints. HA-VLN 2.0 extends far beyond HA-VLN 1.0’s sparse, static settings by introducing unified constraints that substantially increase realism and complexity in both DE and CE: (i) Dynamic Human Models: continuous trajectories from the HAPS 2.0 dataset capturing diverse behaviors and dense crowds; (ii) Personal-Space Enforcement: standardized proximity thresholds (3 m in DE; overlapping radii in CE) to ensure equitable cross-paradigm evaluation; (iii) Human-Focused Instructions: natural-language directives grounded in dynamic human behaviors, requiring precise alignment between text and visual context. All annotations are curated through a validated multi-stage pipeline (Sec. 3), ensuring both realism and reproducibility.
Unified Dynamics and Partial Observability. HA-VLN 2.0 formalizes a unified partially observable Markov decision process (POMDP) spanning both DE and CE settings, whereas HA-VLN 1.0 considered partial observability only in DE. Successor states $s_{t+1}$ depend jointly on agent actions and stochastic human dynamics (e.g., sudden path blockage or unexpected entry). Agents must therefore infer latent human intentions and strategically balance exploration (discovering alternate routes) with exploitation (committing to viable trajectories), reflecting the fundamental trade-offs inherent in navigation through dynamic, human-populated environments.
Key Challenges of DE–CE Synergies. Unifying DE and CE exposes three challenges for socially intelligent navigation: (i) Socially Compliant Navigation: collision-free movement that adapts to evolving personal-space boundaries; (ii) Human-Aligned Instruction Grounding: accurate interpretation of natural-language instructions amid dynamic human activities; (iii) Adaptive Path Re-planning: trajectory adjustment in response to human interactions that modify accessibility. DE supports rapid prototyping and large-scale evaluation, while CE offers motion fidelity indispensable for bridging simulation and real-world deployment. Together, these synergies establish HA-VLN 2.0 as the first benchmark uniting efficient simulation with realistic human-populated environments, motivating a unified simulator and corresponding agent framework introduced next.
Figure 2: HA-VLN Simulator. Unlike HA3D, which modeled sparse and static human activities in discrete settings, HA-VLN incorporates rich and dynamic human behaviors using HAPS 2.0 (172 activities, 486 models, 58k frames). Annotation involves two stages: (i) coarse-to-fine optimization via PSO and multi-view camera setups, and (ii) human-in-the-loop refinement for realistic crowd dynamics. Real-time rendering updates motions through a signaling mechanism, facilitating collision detection and dynamic agent–environment interactions. These improvements bridge discrete evaluation (DE) and realistic continuous navigation (CE), establishing a robust foundation for benchmarks in socially intelligent navigation.
@@ 3 HA-VLN Simulator
To support the unified HA-VLN task, we build a simulator that embeds dynamically moving humans in both discrete and continuous 3D environments. Unlike Li et al. (2024), which treated humans as static obstacles, our simulator models high fidelity motions, interactions among multiple humans, and socially grounded dynamics such as spontaneous movements, group activities, and personal space constraints. Using the upgraded HAPS 2.0 dataset, it improves motion diversity, spatial alignment, and realism over HAPS 1.0 (Table A2) and provides 486 curated sequences across indoor and outdoor scenes. The system includes two modules, HA-VLN-CE (continuous) and HA-VLN-DE (discrete), with a unified API (Sec. 3) for human state queries, dynamic scene updates, and collision checks. Fig. 2 places these components in the agent’s action and observation loop, forming the basis for the annotation, rendering, and interaction mechanisms that follow.
[…]
Human Activity Annotation: Coarse-Level. To integrate HAPS 2.0 into our simulator, we adopt a coarse-to-fine strategy. At the coarse level, each region $\mathbf{R}$ is defined by a label $r$, boundary coordinates $\mathbf{B}_{lo}=(x_{lo},y_{lo},z_{lo})$ and $\mathbf{B}_{hi}=(x_{hi},y_{hi},z_{hi})$, and an object set $\mathbf{O}=\{j_{1},j_{2},\ldots,j_{n}\}$ with positions $\mathbf{p}^{j_{i}}$. We filter $\mathbf{H}$ to retain motions consistent with $r$, forming $\mathbf{H}^{\prime}$. Each motion $h_{i}\in\mathbf{H}^{\prime}$ is paired with an object $j_{i}\in\mathbf{O}$ via semantic similarity, producing $(h_{i},j_{i})$ pairs. Particle Swarm Optimization (PSO) Kennedy & Eberhart (1995) (Alg. A1) then determines the optimal placement $\mathbf{p}_{opt}^{h_{i}}$ around $j_{i}$, bounded by $\mathbf{R}$ and penalized if violating constraints such as maintaining a minimum distance $\epsilon=1$m from other objects or leaving the region (details in Appx. B.2). This yields natural placements that reflect realistic social behaviors and spatial relations.
Human Activity Annotation: Fine-Level. Building on coarse placements, fine-level annotation refinement leverages multi-camera observations, ensuring precise alignment of motions with scene geometry. Inspired by 3D skeleton capture systems Ji et al. (2018); Petrovich et al. (2021), we deploy nine RGB cameras around each human model (Fig. 2; see also Fig. A1). Each camera is located at $\mathbf{p}_{cam}$, shifted by $(\Delta_{x},\Delta_{y},\Delta_{z})$ from the human position $\mathbf{p}_{h}$, with rotation angles $\theta_{lr}$ and $\theta_{ud}$. Horizontal shifts are set as $\Delta_{x},\Delta_{y}=\epsilon$ and the vertical shift as $\Delta_{z}$. For camera $i$ ($i=1,\dots,8$), $\theta_{\text{ud}}^{i}$ is defined as: $\tan\theta_{\text{ud}}^{i}\!\!=\!\!\left\{\begin{array}[]{cl}0&:i\text{~is~odd}\\ \frac{\Delta_{z}}{\sqrt{2}\epsilon}&:i\text{~is~even}\end{array}\right$. and the left-right angle $\theta_{\text{lr}}^{i}\!=\!\frac{\pi i}{8}$, while the overhead camera ($i=9$) has $\theta_{lr}^{9}=0$ and $\theta_{ud}^{9}=\tfrac{\pi}{2}$. This multi-view setup provides dense RGB coverage, enabling fine adjustments to resolve inconsistencies like mesh–object clipping. This stage took over 430 hours of annotation, yielding 529 models across 374 regions in 90 scans.
Human Activity Annotation: Multi-Human Enrichment. In Stage 2 (Fig. 2), we enrich scene diversity and interactions through a human-in-the-loop approach Ding et al. (2024), adding new characters and complex motion paths into regions $\mathbf{R}$ with existing activities $h_{i}$ at positions $\mathbf{p}^{h_{i}}$. Regional context, including objects $\mathbf{O}$ within 6 meters of $h_{i}$ and their positions, is provided to LLMs to generate diverse multi-human scenarios, which are refined in four rounds of manual review for scene consistency. Based on curated descriptions, new motions are placed relative to objects and annotated using the multi-camera method from Stage 1, enabling complex actions such as walking downstairs (details in Appx. B.5). After two annotation stages, the dataset comprises 910 human models across 428 regions in 90 scans (Fig. 3(a)(b)), including 111 outdoor humans, 72 two-person interactions, 59 three-person groups, and 15 four-person groups. Among these, 268 involve complex motions such as climbing stairs, substantially enriching the dataset with realistic behaviors. Detailed statistics are provided in Appx. B.8. This two-stage system enables precise modeling of social interaction spaces and personal boundaries, supporting agents in learning socially appropriate navigation strategies.
@@
Figure 3: Motion Analysis. (a) Top three motions from Stage 1 (without enrichment) and Stage 2 (with enrichment). (b) Overall activity statistics, comparing interaction types, movement distances, and the number of models. Enrichment expands both the variety and dynamic range of human activities.
Real-Time Rendering & Agent Interaction. Beyond static annotation, our simulator continuously renders human motions in real time. A dual-thread producer–consumer architecture (Alg. A2) manages frame updates: Thread 1 enqueues refresh signals, while Thread 2 synchronizes with the agent’s action cycle to process them. Each motion spans up to 120 frames; upon receiving a signal, Thread 2 discards outdated meshes and loads new ones, keeping retrieval delays below 50 ms. Fig. A2 illustrates how multiple humans are simultaneously maintained in a shared environment.
To close the loop, agents perceive these dynamics through a navigation mesh (navmesh) Savva et al. (2019). Collisions are flagged when bounding volumes overlap, i.e., when inter-object distances fall below the sum of their radii, triggering an automatic revert. This integration ensures agents not only experience dynamic and socially realistic environments but also learn to respect personal space and navigate effectively in dense human crowds.
@@
Discrete vs. Continuous Settings. HA-VLN-CE (Continuous) allows agents to move in real-valued increments (e.g., $0.25\,\mathrm{m}$ forward, $15^{\circ}$ turns), supporting fine-grained collision avoidance and adaptive social behavior. As shown in Fig. A4, each scene can host up to 10 humans, with simulation speeds of 30–60 FPS on standard GPUs. HA-VLN-DE (Discrete) extends HA3D Li et al. (2024) by incorporating HAPS 2.0 data across indoor and outdoor environments. Agents hop among panoramic viewpoints while humans move continuously, preserving core social-navigation challenges. To align with continuous motions, we map positions to discrete nodes Li et al. (2024), apply small offsets for refinement, and integrate 627 annotated humans across 90 buildings.
@@
Unified API. We provide a unified API supporting both modes with three core functions: (i) Human State Queries for retrieving bounding volumes, motion frames, and semantic annotations of nearby humans; (ii) Dynamic Scene Updates to notify agents of newly moved humans or environmental changes; and (iii) Collision Checks to evaluate whether a proposed move (e.g., forward step or viewpoint hop) would intersect with a human. By integrating HAPS 2.0, coarse-to-fine annotation, real-time multi-human rendering, and a single API across discrete and continuous settings, the HA-VLN Simulator establishes a comprehensive testbed for socially aware navigation. Figs. A2, A3, and A4 showcase the simulator’s ability to capture diverse human behaviors, while Tables A1 and A2 highlight its advantages over prior simulators and the improvements of HAPS 2.0 relative to HAPS 1.0. Appendix B.7 provides details on environment scales, latency, and usage examples.